2. Income taxes The following table shows income tax rates in Econoland Annual Income Up to $80,000 From $80,001 to $120,000 Over $120,000 Tax Rate 10% 20% 30% This is an example of a income tax. Raphael, a resident of Econoland, currently works 20 hours a week and earns an annual income of $80,000. After paying income taxes, Raphael receives per year. If Raphael works an additional 10 hours a week (30 hours a week total), his annual income will be $120,000. After paying income taxes, Raphael wl receiv per year. In other words, by working an additional 10 hours per week, Raphael will receive an additional
If Raphael works 10 more hours in a week (40 hours a week total), his annual income will further increase to $160,000. After paying income taxes, Raphael will receive what he would receive if he works only 30 hours a week. per year. In other words, by working 10 more hours per week, Raphael will receive more than As Raphael works more, the additional amount he receives from working 10 more hours , which means his cost of leisure
Answer #1

The answers in order of the balnks are as follows:

1. Progressive

Reason: as the tax rate increase with income.

2. $80,000*(1-0.10) = $72,000

Reason: 10% tax is applicable upto $80,000.

3. $72,000+ $40,000*(1-0.2) = 104,000

Thus, additional 104000-72000 = 32,000.

4. 104,000 + 40,000*(1-0.3) = 132,000

Thus, additional 28,000

5. Decreases, decreases

Reason: wage rate is the opportunity cost of leisure.

Thanks!